Session: Mar 2008 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Onsala (Michael Lindqvist at Sun Mar 16 07:10:58 2008) Success :
Higher than normal Tsys in LCP. Holds for other 5cm-exp. also. Not sure why (yet). Otherwise no known problems.

Torun (Kaz Borkowski at Mon Mar 17 08:56:49 2008) Failure : ANT
Not observed due to failure of a resolver in one of the azimuth drives.

Westerbork (Tony Foley at Mon Mar 17 15:04:04 2008) Minor failures :
Single dish was used. Please change the coordinates in the vexfile accordingly. BBC1 was unstable - so Tsys will be hard to calibrate

Effelsberg (Uwe Bach at Tue Mar 18 12:30:37 2008) Minor failures :
Some minutes between 20:29-20:39 UT not on source

Jodrell_Bank (Alastair Gunn at Wed Mar 19 15:06:45 2008) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Wed Mar 19 15:07:10 2008) Success :
No known problems.

Noto (P. Cassaro at Thu Mar 27 11:00:04 2008) Success :